OpenStudy (anonymous):

help please ): Which of the following is not equivalent to the formula d = rt? d over r= t r = dt d over t= r d over r t= 1

OpenStudy (whpalmer4):

\[d = rt\]If we divide both sides by r, we get\[\frac{d}{r} = t\] If we instead divide both sides by t, we get \[\frac{d}{t} = r\]

OpenStudy (anonymous):

so its d over t = r?

OpenStudy (whpalmer4):

No, the formulas that I got were all equivalent.

OpenStudy (whpalmer4):

If you do the same thing to both sides of an equation, it remains true.

OpenStudy (whpalmer4):

One I didn't do: \[d=rt\]Divide both sides by rt \[\frac{d}{rt} = \frac{rt}{rt} = 1\]

OpenStudy (whpalmer4):

You can also do these by looking at the dimensions. t is time. d is length. r is length/time. does r = dt make sense, given that? it would be length * time = length/time, which doesn't work.

OpenStudy (anonymous):

\[\frac{d}{t}=r\] it's like d=rt ;divide t \[\frac{d}{r}=t\] it's like d=rt ;divide r \[\frac{d}{rt}=1\] it's like d=rt ;divide tr \[r=dt\] it's like d=rt but how to get dt multiply t both side \[ dt=rt^{2}\] it should to be dt=r not dt=rt^2 FAILLLLL !!!!

OpenStudy (anonymous):

so how do you know which one is not equivalent?

OpenStudy (whpalmer4):

The one that you couldn't get by valid manipulation. r = dt

OpenStudy (whpalmer4):

My first post showed that A and C were legit. Then I showed that D was legit. That leaves B as the only possibility, and I showed why B wasn't a valid transformation of the original.

OpenStudy (anonymous):

thank you (: